# SAVeD: Semantically Aware Version Discovery

### v1

@@ 5.4. Baseline Comparisons
Since there are currently no existing version discovery methods, we compare SAVeD against a comprehensive set of representative methods for related table and table union search, which we consider to be the closest task. For LLMs, we use the methodology of DiscoverGPT (hu-etal-2025-discovergpt) for comparing version pair candidates. DiscoverGPT builds upon work from DataLore (Lou2024), aiming to be a cross-task methodology utilizing LLMs for database tasks. Intuitively, DiscoverGPT’s should perform well on the task of version discovery, dependent on the LLM engine used for evaluation.
  • Starmie(fan2023semanticsawaredatasetdiscoverydata): State-of-the-art pre-trained RoBERTa-based table understanding model(10.14778/3407790.3407797) adapted for version discovery through supervised fine-tuning
  • DiscoverGPT (Llama2-7B Backbone) (touvron2023llama2openfoundation)- This uses the methodology of DiscoverGPT with the LLama2-7B LLM.
  • DiscoverGPT (Qwen2.5-7B Backbone) :(qwen2025qwen25technicalreport)- This uses the methodology of DiscoverGPT with the Qwen2.5-7B LLM.
  • DiscoverGPT (Gemma2-7B Backbone) : (gemmateam2024gemma2improvingopen)-This uses the methodology of DiscoverGPT with the Gemma2-7B LLM.
@@ 5.5. Design Details
All experiments use a fixed random seed with deterministic CUDA (cuda) operations enabled to ensure reproducible results across runs. Our design features comprehensive seed setting for Python’s random module (van1995python), NumPy(harris2020array), and PyTorch’s CPU and GPU random number generators (DBLP:journals/corr/abs-1912-01703). All experiments were run on an NVIDIA RTX 3090 using CUDA version 12.9. Flash Attention 2 (dao2023flashattention2fasterattentionbetter) was used for faster inference where possible.

### v2

@@ 5.4. Baseline Comparisons
Since there are currently no existing version discovery methods, we compare SAVeD against a comprehensive set of representative methods for related table and table union search, which we consider to be the closest task. For LLMs, we use the methodology of DiscoverGPT (Hu et al., 2025) for comparing version pair candidates. DiscoverGPT builds upon work from DataLore (Lou et al., 2024), aiming to be a cross-task methodology utilizing LLMs for database tasks. Intuitively, DiscoverGPT’s should perform well on the task of version discovery, dependent on the LLM engine used for evaluation.
  • Starmie(Fan et al., 2023): State-of-the-art pre-trained RoBERTa-based table understanding model(Freitag et al., 2020) adapted for version discovery through supervised fine-tuning
  • DiscoverGPT (Llama2-7B Backbone) (Touvron et al., 2023)- This uses the methodology of DiscoverGPT with the LLama2-7B LLM.
  • DiscoverGPT (Qwen2.5-7B Backbone) :(Qwen et al., 2025)- This uses the methodology of DiscoverGPT with the Qwen2.5-7B LLM.
  • DiscoverGPT (Gemma2-7B Backbone) : (Team et al., 2024)-This uses the methodology of DiscoverGPT with the Gemma2-7B LLM.
@@ 5.5. Design Details
All experiments use a fixed random seed with deterministic CUDA (NVIDIA et al., 2020) operations enabled to ensure reproducible results across runs. Our design features comprehensive seed setting for Python’s random module (Van Rossum, 2020), NumPy(Harris et al., 2020), and PyTorch’s CPU and GPU random number generators (Paszke et al., 2019). All experiments were run on an NVIDIA RTX 3090 using CUDA version 12.9. Flash Attention 2 (Dao, 2023) was used for faster inference where possible.